OpenStudy (anonymous):

Solve: negative 1 over 2 x + 10 = −x + 12

OpenStudy (campbell_st):

so the question is \[\frac{-1}{2x + 10} = -x + 12\] or \[\frac{-1}{2x} + 10 = -x + 12\]

OpenStudy (anonymous):

yes but i figured out what the value of "x" is to make this correct its 4

OpenStudy (anonymous):

but thanks for the help